Please help me find the slope-intercept form of the line passing through the point (9,-7) and parallel to the line 7x-9y=7
A.y=7/9x-14
B.y=-7/9x-14
C.y=7/9x+1/14
D.y=9/7x+1/14

Answer by stanbon(75887)

Please help me find the slope-intercept form of the line passing through the point (9,-7) and parallel to the line 7x-9y=7
-----------
The given line has slope = 7/9
----
Any parallel line must have slope = 7/9
-----
Form: y = mx + b
-7 = (7/9)(9) + b
b = -14
==========
Equation:
y = (7/9)x -14
=====================
Cheers,
Stan H.
=====================

A.y=7/9x-14
B.y=-7/9x-14
C.y=7/9x+1/14
D.y=9/7x+1/14